# Discovering Interpretable Programmatic Policies via Multimodal LLM-assisted Evolutionary Search

### v1

@@ E.1 Templates for Constructing Prompts for Specific Operators
This section details the prompt templates utilized for policy generation within the MLES framework. Figures 10, 11, 12, and 13 illustrate the specific prompt templates for operators E1, E2, M1_M, and M2_M, respectively. In all presented templates, black text indicates fixed components, red placeholders denote task-specific elements, and blue placeholders represent variables that evolve throughout the evolutionary process.
@@ 3. Thirdly, implement the following Python function:
Execution results visualization for the algorithm: “[Interpretabal Behavioral Evidence Placeholder]”
@@ 3. Implement the enhanced algorithm using the following Python function template:
Execution results visualization for the algorithm: “[Interpretable Behavioral Evidence Placeholder]”
@@ 1. Parameter Analysis:
 - Identify all key parameters and their functions.
 - Determine which parameters should be modified to improve results.
 - Explain why these specific changes would help.
@@ E.2 Task-Specific Prompts Used in Experiments
This section demonstrates the design of task descriptions and code templates for applying MLES, using two control problems from our experiments as examples. Briefly, the task description communicates the requirements for the control policy to be designed, while the code template defines the programmatic policy’s code interface that MLES needs to generate. Figures 14 and 15 present the task descriptions provided to MLES for the Lunar Lander and Car Racing tasks, respectively. Figures 16 and 17 show the corresponding code templates.
@@
Within the function declaration and docstring, we must pre-define the parameters that the policy and environment will exchange, ensuring compatibility with the environment’s interface. This involves: (1) identifying the parameters the environment or system provide to the policy and defining them as input variables in the function declaration; (2) detailing the type, dimension, range, and meaning of these input parameters in the docstring; and (3) clearly specifying the policy’s output in the docstring, typically the actions the agent can take. Additionally, we can incorporate prior knowledge and hints within the docstring to enhance the effectiveness of policies generated by MLES. As shown in Figure 17, we can include ”Notes” sections within the docstring to help MLES better understand the environment and offer specific guidance. For the function body, we can simply provide a return statement to ensure function completeness, or we can additionally include existing expert heuristics. Generally, providing some expert knowledge facilitates the policy discovery process of MLES.
1import numpy as np
2
3def choose_action(s: list, last_action: int, s_pre: list) -> int:
4 """
5 Selects an action for the Lunar Lander to achieve a safe landing at the target location (0, 0).
6 Args:
7 s (list or np.ndarray): The current state of the lander. Elements:
8 s[0] - horizontal position (x)
9 s[1] - vertical position (y)
10 s[2] - horizontal velocity (v_x)
11 s[3] - vertical velocity (v_y)
12 s[4] - angle (radians)
13 s[5] - angular velocity
14 s[6] - 1 if the first leg is in contact with the ground, else 0
15 s[7] - 1 if the second leg is in contact with the ground, else 0
16 last_action (int): The action taken in the previous step. One of:
17 0 - do nothing
18 1 - fire left orientation engine
19 2 - fire main (upward) engine
20 3 - fire right orientation engine
21 s_pre (list or np.ndarray): The state of the lander *before* the last action was executed.
22 Returns:
23 int: The chosen action for the next step. One of:
24 0 - do nothing
25 1 - fire left orientation engine
26 2 - fire main (upward) engine
27 3 - fire right orientation engine
28 """
29 angle_targ = s[0] * 0.5 + s[2] * 1.0 # angle should point towards center
30 if angle_targ > 0.4:
31 angle_targ = 0.4 # more than 0.4 radians (22 degrees) is bad
32 if angle_targ < -0.4:
33 angle_targ = -0.4
34 hover_targ = 0.55 * np.abs(
35 s[0]
36 ) # target y should be proportional to horizontal offset
37
38 angle_todo = (angle_targ - s[4]) * 0.5 - (s[5]) * 1.0
39 hover_todo = (hover_targ - s[1]) * 0.5 - (s[3]) * 0.5
40
41 if s[6] or s[7]: # legs have contact
42 angle_todo = 0
43 hover_todo = (
44 -(s[3]) * 0.5
45 ) # override to reduce fall speed, that’s all we need after contact
46
47 a = 0
48 if hover_todo > np.abs(angle_todo) and hover_todo > 0.05:
49 a = 2
50 elif angle_todo < -0.05:
51 a = 3
52 elif angle_todo > +0.05:
53 a = 1
54 return a
@@ 22 Returns:
1import numpy as np
2import cv2
3def choose_action(observation, car_speed, pre_action, pre_observation):
4 """
5 Determine the next action for the Car Racing agent.
6 This function takes into account the current state (observation and speed), the previous action, and the previous observation.
7 Notes:
8 - The car in this environment is a powerful rear-wheel-drive vehicle. Avoid accelerating while turning sharply,
9 as this can easily lead to loss of control.
10 - Occasionally, track segments (e.g., after a U-turn) may appear in the observation but are not part of the immediate drivable path. These should be distinguished to avoid premature or incorrect decisions.
11 - Avoid coming to a complete stop, as this may prevent the car from finishing the race.
12 Args:
13 observation (np.ndarray): The current state observed by the agent, represented as a 96x96 RGB image of the car and race track from a top-down view (shape: (96, 96, 3)).
14 car_speed (float): The current speed of the car.
15 pre_action (np.ndarray): The action taken by the car in the previous step, represented as a 3-element array.
16 pre_observation (np.ndarray): The observation received when the previous action was taken. It has the same shape and format as ‘observation‘ (i.e., a 96x96 RGB image).
17 Returns:
18 np.ndarray: The action selected by the agent for the next step, represented as an array of shape (3,) where:
19 - Index 0: Steering, where -1 is full left, +1 is full right (range: [-1, 1]).
20 - Index 1: Gas, (range: [0, 1]).
21 - Index 2: Braking, (range: [0, 1]).
22 """
23 action = np.array([0.0, 0.0, 0.0])
24 # Gray track detection parameters (RGB 95-115 range with +-5% tolerance)
25 gray_low = 95
26 gray_high = 115
27
28 # Create 3D gray detection mask (all RGB channels within range)
29 gray_mask = (
30 (observation[:, :, 0] >= gray_low) & (observation[:, :, 0] <= gray_high) &
31 (observation[:, :, 1] >= gray_low) & (observation[:, :, 1] <= gray_high) &
32 (observation[:, :, 2] >= gray_low) & (observation[:, :, 2] <= gray_high)
33 )
34
35 gray_indices = np.argwhere(gray_mask)
36 center_x = np.mean(gray_indices[:, 1]) if len(gray_indices) > 0 else observation.shape[1] // 2
37 car_position = observation.shape[1] // 2
38 offset = center_x - car_position
39
40 steering_angle = np.clip(offset / 100.0, -1.0, 1.0)
41 action[0] = steering_angle
42
43 if abs(offset) > 10:
44 action[1] = 0.0
45 action[2] = 0.2
46 else:
47 action[1] = 0.8
48 action[2] = 0.0
49
50 gray_density = np.sum(gray_mask) / (observation.shape[0] * observation.shape[1])
51 if gray_density < 0.1:
52 action[1] = 0.4
53 action[2] = 0.3
54 return action
@@ Appendix F Generated Programmatic Policy
Fig. 18 shows the code for the best policy discovered for the Lunar Lander task. The thought of this policy provided by MLES is: ”The core idea of the new algorithm is to further increase the sensitivity to vertical velocity, making the lander more responsive to abrupt changes, while also lowering the action thresholds to enable quicker decisions for stability control.” Fig.19 displays the performance of this policy across several instances, highlighting its ability to achieve safe landings consistently.
1def choose_action(s: list, last_action: int, s_pre: list) -> int:
2 """
3 Selects an action for the Lunar Lander to achieve a safe landing at the target location (0, 0).
4 Args:
5 s (list or np.ndarray): The current state of the lander. Elements:
6 s[0] - horizontal position (x)
7 s[1] - vertical position (y)
8 s[2] - horizontal velocity (v_x)
9 s[3] - vertical velocity (v_y)
10 s[4] - angle (radians)
11 s[5] - angular velocity
12 s[6] - 1 if the first leg is in contact with the ground, else 0
13 s[7] - 1 if the second leg is in contact with the ground, else 0
14 last_action (int): The action taken in the previous step. One of:
15 0 - do nothing
16 1 - fire left orientation engine
17 2 - fire main (upward) engine
18 3 - fire right orientation engine
19 s_pre (list or np.ndarray): The state of the lander *before* the last action was executed.
20 Returns:
21 int: The chosen action for the next step. One of:
22 0 - do nothing
23 1 - fire left orientation engine
24 2 - fire main (upward) engine
25 3 - fire right orientation engine
26 """
27 angle_targ = (s[0] * 0.5 + s[2] * 0.5) # Designed for better orientation control
28 angle_targ = np.clip(angle_targ, -0.6, 0.6) # Slightly wider bounds for orientation
29
30 # Update hover target with amplified sensitivity to vertical speed
31 hover_targ = 0.3 * np.abs(s[0]) + 0.7 * (s[3] ** 2) # Greater weight for hover stabilizing
32
33 # Calculate actions for adjustments
34 angle_todo = (angle_targ - s[4]) * 1.6 - (s[5]) * 1.2 # More aggressive response for angle adjustments
35 hover_todo = (hover_targ - s[1]) * 1.3 - (s[3]) * 0.4 # Enhanced weight for vertical control
36
37 # Stabilizing when legs are in contact
38 if s[6] or s[7]:
39 angle_todo = 0 # Maintain upright position priority
40 hover_todo -= (s[3]) * 0.4 # Stronger adjustment for vertical speed reduction
41
42 # Decision-making based on refined thresholds and more responsive measures
43 a = 0
44 if hover_todo > np.abs(angle_todo) and hover_todo > 0.10: # Slightly tighter hover threshold
45 a = 2 # Fire main engine
46 elif angle_todo < -0.15: # Increased sensitivity for left engine
47 a = 3
48 elif angle_todo > 0.15: # Increased sensitivity for right engine
49 a = 1
50 return a
@@ 20 Returns:
1def choose_action(observation, car_speed, pre_action, pre_observation):
2 """
3 Determine the next action for the Car Racing agent.
4 Args:
5 observation (np.ndarray): The current state observed by the agent, represented as a 96x96 RGB image of the car and race track from a top-down view (shape: (96, 96, 3)).
6 car_speed (float): The current speed of the car.
7 pre_action (np.ndarray): Action taken by the agent in the previous step, represented as a 3-element array.
8 pre_observation (np.ndarray): The observation received when the previous action was taken. It has the same shape and format as ‘observation‘ (i.e., a 96x96 RGB image).
9 Returns:
10 np.ndarray: The action selected by the agent for the next step, represented as an array of shape (3,) where:
11 - Index 0: Steering, where -1 is full left, +1 is full right (range: [-1, 1]).
12 - Index 1: Gas, (range: [0, 1]).
13 - Index 2: Braking, (range: [0, 1]).
14 """
15 hsv = cv2.cvtColor(observation, cv2.COLOR_RGB2HSV)
16 # Define masks for track, off-track, and curbs detection
17 track_mask = cv2.inRange(hsv, (0, 0, 60), (179, 255, 200)) # Track pixels
18 off_track_mask = cv2.inRange(hsv, (40, 100, 100), (90, 255, 255)) # Off-track pixels
19 curbs_mask = cv2.inRange(hsv, (180, 100, 100), (255, 255, 255)) # Curbs detection
20
21 # Analyze the center region for better decision making
22 height, width = track_mask.shape
23 center_region = track_mask[int(height / 3):int(2 * height / 3), int(width / 4):int(3 * width / 4)]
24 # Initialize the action array
25 action = np.zeros(3)
26 # Identify whether track is present and where
27 track_detected = np.sum(center_region) > 0
28 if not track_detected:
29 action[2] = 1.0 # Full brake if no track detected
30 else:
31 # Identify track pixel positions
32 track_pixels = np.column_stack(np.where(center_region > 0))
33 mean_x = np.mean(track_pixels[:, 1]) if len(track_pixels) > 0 else center_region.shape[1] / 2
34
35 # Steering adjustment with consideration for immediate curvature
36 steering = (mean_x - (center_region.shape[1] / 2)) / (center_region.shape[1] / 2)
37 action[0] = np.clip(steering, -1, 1) # Normalize steering value
38
39 # Adaptive throttle management based on speed
40 if car_speed < 1.0:
41 action[1] = 1.0 # Full throttle if the car is very slow
42 else:
43 # Adjust throttle considering steering
44 if abs(action[0]) > 0.5:
45 action[1] = max(0.0, 1.0 - abs(action[0])) # Reduce throttle during significant turns
46 else:
47 action[1] = min(1.0, action[1] + 0.05) # Gradual increase on straights
48
49 # Enhanced braking logic considering sharp turns
50 if abs(action[0]) > 0.6 and car_speed > 2.0:
51 action[2] = min(0.4, action[2] + 0.1) # Moderate braking for sudden turns
52 else:
53 action[2] = 0.0 # No brake needed on straights
54
55 # Prevent the car from stalling
56 if np.all(pre_action == [0, 0, 1]) and car_speed < 0.5:
57 action[1] = max(action[1], 0.5) # Ensure some throttle to avoid stalling
58 return action
@@ 9 Returns:
Figure 21: Performance of the best discovered policy on different instances in the Car Racing task. These figure shows the car’s smooth track completion and race line strategy, similar to human racing behavior.

### v2

@@ E.1 Templates for Constructing Prompts for Specific Operators
This section details the prompt templates utilized for policy generation within the MLES framework. Figs. 11, 12, 13, and 14 illustrate the specific prompt templates for operators E1, E2, M1_M, and M2_M, respectively. In all presented templates, black text indicates fixed components, red placeholders denote task-specific elements, and blue placeholders represent variables that evolve throughout the evolutionary process.
@@ 3. Thirdly, implement the following Python function:
Execution results visualization for the algorithm: “[Behavioral Evidence Placeholder]”
@@ 3. Implement the enhanced algorithm using the following Python function template:
Execution results visualization for the algorithm: “[Behavioral Evidence Placeholder]”
@@ 1. Parameter Analysis:
- Identify all key parameters and their functions.
- Determine which parameters should be modified to improve results.
- Explain why these specific changes would help.
@@ E.2 Task-Specific Prompts Used in Experiments
This section demonstrates the design of task descriptions and code templates for applying MLES, using two control problems from our experiments as examples. Briefly, the task description communicates the requirements for the control policy to be designed, while the code template defines the programmatic policy’s code interface that MLES needs to generate. Figs. 15 and 16 present the task descriptions provided to MLES for the Lunar Lander and Car Racing tasks, respectively. Figs. 17 and 18 show the corresponding code templates.
@@
Within the function declaration and docstring, we must pre-define the parameters that the policy and environment will exchange, ensuring compatibility with the environment’s interface. This involves: (1) identifying the parameters the environment or system provide to the policy and defining them as input variables in the function declaration; (2) detailing the type, dimension, range, and meaning of these input parameters in the docstring; and (3) clearly specifying the policy’s output in the docstring, typically the actions the agent can take. Additionally, we can incorporate prior knowledge and hints within the docstring to enhance the effectiveness of policies generated by MLES. As shown in Fig. 18, we can include “Notes” sections within the docstring to help MLES better understand the environment and offer specific guidance. For the function body, we can simply provide a return statement to ensure function completeness, or we can additionally include existing expert heuristics. Generally, providing some expert knowledge facilitates the policy discovery process of MLES.
import numpy as np
def choose_action(s: list, last_action: int, s_pre: list) -> int:
Selects an action for the Lunar Lander to achieve a safe landing at the target location (0, 0).
Args:
s (list or np.ndarray): The current state of the lander. Elements:
s[0] - horizontal position (x)
s[1] - vertical position (y)
s[2] - horizontal velocity (v_x)
s[3] - vertical velocity (v_y)
s[4] - angle (radians)
s[5] - angular velocity
s[6] - 1 if the first leg is in contact with the ground, else 0
s[7] - 1 if the second leg is in contact with the ground, else 0
last_action (int): The action taken in the previous step. One of:
0 - do nothing
1 - fire left orientation engine
2 - fire main (upward) engine
3 - fire right orientation engine
s_pre (list or np.ndarray): The state of the lander *before* the last action was executed.
Returns:
int: The chosen action for the next step. One of:
0 - do nothing
1 - fire left orientation engine
2 - fire main (upward) engine
3 - fire right orientation engine
angle_targ = s[0] * 0.5 + s[2] * 1.0 # angle should point towards center
if angle_targ > 0.4:
angle_targ = 0.4 # more than 0.4 radians (22 degrees) is bad
if angle_targ < -0.4:
angle_targ = -0.4
hover_targ = 0.55 * np.abs(
s[0]
) # target y should be proportional to horizontal offset
angle_todo = (angle_targ - s[4]) * 0.5 - (s[5]) * 1.0
hover_todo = (hover_targ - s[1]) * 0.5 - (s[3]) * 0.5
if s[6] or s[7]: # legs have contact
angle_todo = 0
hover_todo = (
-(s[3]) * 0.5
) # override to reduce fall speed, that’s all we need after contact
a = 0
if hover_todo > np.abs(angle_todo) and hover_todo > 0.05:
a = 2
elif angle_todo < -0.05:
a = 3
elif angle_todo > +0.05:
a = 1
return a
@@
import numpy as np
import cv2
def choose_action(observation, car_speed, pre_action, pre_observation):
Determine the next action for the Car Racing agent.
This function takes into account the current state (observation and speed), the previous action, and the previous observation.
Notes:
- The car in this environment is a powerful rear-wheel-drive vehicle. Avoid accelerating while turning sharply,
as this can easily lead to loss of control.
- Occasionally, track segments (e.g., after a U-turn) may appear in the observation but are not part of the immediate drivable path. These should be distinguished to avoid premature or incorrect decisions.
- Avoid coming to a complete stop, as this may prevent the car from finishing the race.
Args:
observation (np.ndarray): The current state observed by the agent, represented as a 96x96 RGB image of the car and race track from a top-down view (shape: (96, 96, 3)).
car_speed (float): The current speed of the car.
pre_action (np.ndarray): The action taken by the car in the previous step, represented as a 3-element array.
pre_observation (np.ndarray): The observation received when the previous action was taken. It has the same shape and format as ‘observation‘ (i.e., a 96x96 RGB image).
Returns:
np.ndarray: The action selected by the agent for the next step, represented as an array of shape (3,) where:
- Index 0: Steering, where -1 is full left, +1 is full right (range: [-1, 1]).
- Index 1: Gas, (range: [0, 1]).
- Index 2: Braking, (range: [0, 1]).
action = np.array([0.0, 0.0, 0.0])
# Gray track detection parameters (RGB 95-115 range with +-5% tolerance)
gray_low = 95
gray_high = 115
# Create 3D gray detection mask (all RGB channels within range)
gray_mask = (
(observation[:, :, 0] >= gray_low) & (observation[:, :, 0] <= gray_high) &
(observation[:, :, 1] >= gray_low) & (observation[:, :, 1] <= gray_high) &
(observation[:, :, 2] >= gray_low) & (observation[:, :, 2] <= gray_high)
gray_indices = np.argwhere(gray_mask)
center_x = np.mean(gray_indices[:, 1]) if len(gray_indices) > 0 else observation.shape[1] // 2
car_position = observation.shape[1] // 2
offset = center_x - car_position
steering_angle = np.clip(offset / 100.0, -1.0, 1.0)
action[0] = steering_angle
if abs(offset) > 10:
action[1] = 0.0
action[2] = 0.2
else:
action[1] = 0.8
action[2] = 0.0
gray_density = np.sum(gray_mask) / (observation.shape[0] * observation.shape[1])
if gray_density < 0.1:
action[1] = 0.4
action[2] = 0.3
return action
@@ Appendix F Generated Programmatic Policy
Fig. 19 shows the code for the best policy discovered for the Lunar Lander task. The thought of this policy provided by MLES is: “The core idea of the new algorithm is to further increase the sensitivity to vertical velocity, making the lander more responsive to abrupt changes, while also lowering the action thresholds to enable quicker decisions for stability control.” Fig.20 displays the performance of this policy across several instances, highlighting its ability to achieve safe landings consistently.
def choose_action(s: list, last_action: int, s_pre: list) -> int:
Selects an action for the Lunar Lander to achieve a safe landing at the target location (0, 0).
Args:
s (list or np.ndarray): The current state of the lander. Elements:
s[0] - horizontal position (x)
s[1] - vertical position (y)
s[2] - horizontal velocity (v_x)
s[3] - vertical velocity (v_y)
s[4] - angle (radians)
s[5] - angular velocity
s[6] - 1 if the first leg is in contact with the ground, else 0
s[7] - 1 if the second leg is in contact with the ground, else 0
last_action (int): The action taken in the previous step. One of:
0 - do nothing
1 - fire left orientation engine
2 - fire main (upward) engine
3 - fire right orientation engine
s_pre (list or np.ndarray): The state of the lander *before* the last action was executed.
Returns:
int: The chosen action for the next step. One of:
0 - do nothing
1 - fire left orientation engine
2 - fire main (upward) engine
3 - fire right orientation engine
angle_targ = (s[0] * 0.5 + s[2] * 0.5) # Designed for better orientation control
angle_targ = np.clip(angle_targ, -0.6, 0.6) # Slightly wider bounds for orientation
# Update hover target with amplified sensitivity to vertical speed
hover_targ = 0.3 * np.abs(s[0]) + 0.7 * (s[3] ** 2) # Greater weight for hover stabilizing
# Calculate actions for adjustments
angle_todo = (angle_targ - s[4]) * 1.6 - (s[5]) * 1.2 # More aggressive response for angle adjustments
hover_todo = (hover_targ - s[1]) * 1.3 - (s[3]) * 0.4 # Enhanced weight for vertical control
# Stabilizing when legs are in contact
if s[6] or s[7]:
angle_todo = 0 # Maintain upright position priority
hover_todo -= (s[3]) * 0.4 # Stronger adjustment for vertical speed reduction
# Decision-making based on refined thresholds and more responsive measures
a = 0
if hover_todo > np.abs(angle_todo) and hover_todo > 0.10: # Slightly tighter hover threshold
a = 2 # Fire main engine
elif angle_todo < -0.15: # Increased sensitivity for left engine
a = 3
elif angle_todo > 0.15: # Increased sensitivity for right engine
a = 1
return a
@@
def choose_action(observation, car_speed, pre_action, pre_observation):
Determine the next action for the Car Racing agent.
Args:
observation (np.ndarray): The current state observed by the agent, represented as a 96x96 RGB image of the car and race track from a top-down view (shape: (96, 96, 3)).
car_speed (float): The current speed of the car.
pre_action (np.ndarray): Action taken by the agent in the previous step, represented as a 3-element array.
pre_observation (np.ndarray): The observation received when the previous action was taken. It has the same shape and format as ‘observation‘ (i.e., a 96x96 RGB image).
Returns:
np.ndarray: The action selected by the agent for the next step, represented as an array of shape (3,) where:
- Index 0: Steering, where -1 is full left, +1 is full right (range: [-1, 1]).
- Index 1: Gas, (range: [0, 1]).
- Index 2: Braking, (range: [0, 1]).
hsv = cv2.cvtColor(observation, cv2.COLOR_RGB2HSV)
# Define masks for track, off-track, and curbs detection
track_mask = cv2.inRange(hsv, (0, 0, 60), (179, 255, 200)) # Track pixels
off_track_mask = cv2.inRange(hsv, (40, 100, 100), (90, 255, 255)) # Off-track pixels
curbs_mask = cv2.inRange(hsv, (180, 100, 100), (255, 255, 255)) # Curbs detection
# Analyze the center region for better decision making
height, width = track_mask.shape
center_region = track_mask[int(height / 3):int(2 * height / 3), int(width / 4):int(3 * width / 4)]
# Initialize the action array
action = np.zeros(3)
# Identify whether track is present and where
track_detected = np.sum(center_region) > 0
if not track_detected:
action[2] = 1.0 # Full brake if no track detected
else:
# Identify track pixel positions
track_pixels = np.column_stack(np.where(center_region > 0))
mean_x = np.mean(track_pixels[:, 1]) if len(track_pixels) > 0 else center_region.shape[1] / 2
# Steering adjustment with consideration for immediate curvature
steering = (mean_x - (center_region.shape[1] / 2)) / (center_region.shape[1] / 2)
action[0] = np.clip(steering, -1, 1) # Normalize steering value
# Adaptive throttle management based on speed
if car_speed < 1.0:
action[1] = 1.0 # Full throttle if the car is very slow
else:
# Adjust throttle considering steering
if abs(action[0]) > 0.5:
action[1] = max(0.0, 1.0 - abs(action[0])) # Reduce throttle during significant turns
else:
action[1] = min(1.0, action[1] + 0.05) # Gradual increase on straights
# Enhanced braking logic considering sharp turns
if abs(action[0]) > 0.6 and car_speed > 2.0:
action[2] = min(0.4, action[2] + 0.1) # Moderate braking for sudden turns
else:
action[2] = 0.0 # No brake needed on straights
# Prevent the car from stalling
if np.all(pre_action == [0, 0, 1]) and car_speed < 0.5:
action[1] = max(action[1], 0.5) # Ensure some throttle to avoid stalling
return action
@@
Figure 22: Performance of the best discovered policy on different instances in the Car Racing task. These figures show the car’s smooth track completion and race line strategy, similar to human racing behavior.
Appendix G The Use of Large Language Models
LLMs are utilized in two primary ways in this study: first, to serve as programmatic policy generators within the MLES framework; second, to help refine the manuscript by enhancing the clarity of the language, while the authors maintain full responsibility for the content and direction.